OpenStudy (anonymous):

solve this in substitution way 3x+2y=9 and -2y+y=4

OpenStudy (anonymous):

have you solved the previous problem?

OpenStudy (anonymous):

yes.it actually work but finding LCD

OpenStudy (anonymous):

what LCD?

OpenStudy (anonymous):

-2y+y=4 -y=4 y=-4 3x+2(-4)=9 3x-8=9 3x=1 x=1/3

OpenStudy (anonymous):

OUR teacher said that we have to find the LCD (least common denominator )

OpenStudy (anonymous):

Actually, you don't need to find any LCDs because first: there are no fractions to simplify then second, you can simply get the value of one variable and substitute it to the other. It's that simple.

OpenStudy (anonymous):

okey.thanks